Key Assessment PAT 6: Sayles Chapter 14 Application Exercise 2

Beginning January 1, Community Hospital had 150 inpatient beds. On July 1, the bed count was increased to 200. From January 1 to June 30th, (non-leap year) the hospital provided 26,106 inpatient service days and from July 1 to December 31, there were 35,120 service days. There were 59,788 total discharge days. In addition, there were 9,601 admissions and 9,588 discharges.

Answer the following questions below

1.)  What is the average daily census for the year? Rounded to a whole number

2.)What is the bed occupancy rate for the following? Rounded to two decimal places

January – June:

3.) What is the bed occupancy rate for the following? Round to two decimal places

July – December:

4) What is the bed occupancy rate for the following? Round to two decimal places

5.)  What is the ALOS for the year? Round to one decimal place
Answer #1

1)

Daily inpatient census = Total number of inpatient service days / Total number of days = (26106+35120)/365 = 167.74

2)

Bed occupancy rate = 100*average daily census / number of beds = 100*(26106/181)/150 = 96.15%

3)

Bed occupancy rate = 100*(35120/184)/200 = 95.43%

4)

The annual bed occupancy rate is 96.15*181/365 + 95.43*184/365 = 95.78

5)

The average length of stay = total discharge days/number of discharges = 59788/9588 = 6.23 days